In a particular stress relieving process, a vessel which was 3/4 filled with water at 19 degrees celcius was then completely filled with water in such a way that the final temperature of the water in the vessel doubled. Calculate the temperature of the water that was added.
1
Expert's answer
2017-12-15T09:51:06-0500
From the thermal equilibrium:
3/4 mc(2T-T)=1/4 mc(T'-T)
3/4 (T)=1/4 (T'-T)
(T'-T)=3T
T^'=4T=4(19)=76℃